# Calibrating Bayesian Tension Statistics using Neural Ratio Estimators

### v1

@@ I Introduction
Independently confirming conclusions about the nature of our Universe from one experiment with another is crucial to the advance of knowledge. When the inference from two different experiments disagree with each other, we call this tension. Tension between different datasets raises questions about the need for new physics and better descriptions of our instruments and systematics. The $H_{0}$ and $\sigma_{8}$ tensions (e.g. Knox and Millea, 2020; Efstathiou, 2020; Amon and Efstathiou, 2022; Preston et al., 2023; Dark Energy Survey and Kilo-Degree Survey Collaboration, 2023) are the most commonly encountered examples in cosmology, but other examples include observations of the 21-cm signal from cosmic dawn (Singh et al., 2022a), tensions in the amplitude of the matter power spectrum (Battye et al., 2015) and tension in estimates of the curvature of the Universe Handley (2021). A historical example of tension is in the measurement of the matter density $\Omega_{m}$ Peebles (1984); Krauss and Turner (1995); Ostriker and Steinhardt (1995); Efstathiou et al. (1990) which was resolved by the discovery of the accelerating universe Riess et al. (1998). A detailed review of cosmological tensions can be found in Abdalla et al. (2022).
@@
Many different measures of tension have been proposed and are widely used in cosmological studies. Examples include Bayesian Suspiciousness Handley and Lemos (2019), estimators of the probability of observed parameter differences Raveri et al. (2020); Raveri and Doux (2021), Goodness of fit degradation Raveri and Hu (2019) and Eigentension Park and Rozo (2020). These tension statistics are summarised and reviewed in Charnock et al. (2017); DES Collaboration (2021) and Saraivanov et al. (2024). It is common practice to rephrase these tension metrics in $\sigma$ units of tension, corresponding to probabilities on a one dimensional normal distribution. When expressed in this way, one would expect that the various tension metrics predict the same level of tension or concordance between different datasets. However, this is often not the case due to the various assumptions that are made when defining the statistics. To tackle this issue, we can try to define tension metrics that do not make these assumptions, however, the tension statistics often lose some of their interpretability when we do this. Instead, we try to calibrate out these assumptions in sensible ways. Calibration of tension statistics is an important and often overlooked step that needs to be taken to correctly interrogate the tension between different datasets.
@@
A commonly used metric of tension is $R$ corresponding to the ratio of a joint evidence and the product of individual evidences for two different datasets under a common model. The $R$ statistic was first proposed in (Marshall et al., 2006) and has been used to quantify tension in a number of cosmological studies (e.g. Trotta, 2008; Seehars et al., 2016). $R$ has also been used to perform model comparison in some works, although the authors of Cortês and Liddle (2024) showed that this approach to model comparison is incomplete.
@@
The ratio $R$ suffers from a non-trivial dependence on the prior that is not always accounted for properly (Handley and Lemos, 2019). In Handley and Lemos (2019) the authors showed that as one decreases the prior width on the common parameters in the model of the two datasets, then the tension between the datasets should increase and $R$ should decrease. Intuitively, one can see that it is more satisfying if the two experiments favour parameters that are close together given a wide prior in comparison to a narrow prior. However, what constitutes ‘narrow’ and ‘wide’ is problem specific and subjective, making the interpretation of $R$ difficult. We would like to calibrate out the prior dependence.
The authors of Handley and Lemos (2019) propose an alternative statistic that is closely related to $R$ called the Suspiciousness $S$ which is insensitive to the prior provided the change in prior do not impact the posterior significantly. In Cuceu et al. (2019) the authors convert $S$ into $\sigma$s of tension, however, this requires an estimate of the number of constrained dimensions $d$ in the joint analysis. To estimate $d$ they use the Bayesian (sometimes referred to as Gaussian) Model Dimensionality, however, this is a poor estimator of $d$ if the posterior is significantly non-Gaussian, as is often the case in cosmology.
@@
In DES Collaboration (2021) the authors demonstrate that simulations can be used to calibrate tension metrics with the Planck data and Dark Energy Survey (DES) data. They proposed taking a fiducial set of parameters, such as the maximum posterior point for Planck, shifting these parameter values by some posterior-informed step sizes to induce a known degree of tension, simulate the now in tension DES observations and calculate the value of ones chosen tension metric between the real Planck data and the simulation. To do this, however, one often has to run expensive sampling algorithms on the simulated data to calculate tension statistics such as $R$ and $S$.
@@
We propose calibrating the prior dependence of $R$ using neural ratio estimation (NRE) (e.g. Cranmer et al., 2020; Miller et al., 2021; Cole et al., 2022). NREs are classifiers, with interpretable outputs, that determine whether two quantities are drawn from independent distributions or a joint distribution. We show that the output of an NRE trained on simulations of two experiments observables can be interpreted as $R$ and that an appropriately trained NRE can be used to calibrate for the prior dependence of the dimensionless $R$ statistic. Since $R$ requires the calculation of three Bayesian evidences, it is an expensive statistic to evaluate using traditional methods like nested sampling. We show that $R$ can be accessed with a significantly smaller computational overhead using cutting edge machine learning tools. We call our NRE setup the tensionnet.
@@
In section II we summarise Bayesian inference and give more details about $R$. In section III we discuss the interpretation of $R$ and follow this with a discussion on NREs in section IV. We discuss calibrating $R$ with NREs in section V. We then test our method on toy examples with known in concordance $R$ distributions in section VI. We then apply the tensionnet to a toy example inspired by 21-cm cosmology and to assess the tension between Baryon Acoustic Oscillations (BAO) observations from the Dark Energy Spectroscopic Instrument (DESI) and the Sloan Digital Sky Survey (SDSS), in section VII. We consider some limitations of our method in section VIII and conclude in section IX.
@@ II Bayesian Inference and Tension Statistics
[…]
An efficient and accurate way to recover both the evidence and the posterior is with nested sampling (Skilling, 2006; Ashton et al., 2022) although other methods exist (e.g. Trotta, 2007; Heavens et al., 2017; Srinivasan et al., 2024; Foreman-Mackey et al., 2013; Polanska et al., 2023).
@@
[…]
@@
where we have assumed the data sets are independent, and the angled brackets represent averages over the distributions $P_{A}$ and $P_{B}$ (Handley and Lemos, 2019). For a uniform prior, $\pi=1/V$ where $V$ is the volume, one can see from eq. 4 that if the prior is made smaller than $R$ being proportional to $V$ also decreases. This logic generalises to more complicated priors.
@@ III Interpreting $R$
$R$ has the attractive properties of being dimensionally consistent, parameterisation invariant and symmetric Handley and Lemos (2019). It is typically interpreted with respect to a value of 1 with $R\ll 1$ corresponding to inconsistent datasets and $R\gg 1$ to consistent data. However, this interpretation does not tell us the degree to which our datasets are in tension given the prior and model choice. To try and quantify the tension between observations from the Dark Energy Survey and Planck, the authors of Dark Energy Survey Collaboration (2018) interpreted $R$ on a Jefferys’ scale(Jeffreys, 1983). The Jefferys’ scale is, however, somewhat arbitrary.
In Handley and Lemos (2019) the authors showed that
@@
When interpreting $R$ one has to keep in mind the impact which the prior has on its value. Reducing the width of the prior will increase the apparent tension between the datasets by reducing the value of $R$. The authors of Handley and Lemos (2019) suggest repeating our analysis with sensible modifications to the prior distribution to determine how stable the value of $R$ is and consequently our conclusions regarding the tension between different datasets.
@@ IV Neural Ratio Estimation
$D_{A}$$\vdots$ $D_{B}$$\vdots$ $\log R$ $p=S_{\sigma}(\log R)$ Loss Function: $l=\frac{1}{N}\bigg{[}\sum_{i}^{N}y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})\bigg{]}$ Neural Ratio Estimation of $\log R=\log\frac{P(D_{A},D_{B})}{P(D_{A})P(D_{B})}$
Figure 1: A schematic of the neural ratio estimator (NRE) used in this work, which we refer to as a tensionnet. The NRE is trained on matched and mismatched pairs of simulated observations from two different experiments $A$ and $B$ and outputs an estimate of the tension statistic $R$. The network is trained using the binary cross entropy loss function.
@@ V Calibrating $R$ with NREs
We propose that the true observed tension $R_{\mathrm{obs}}$ is calculated using nested sampling (e.g. Handley and Lemos, 2019) or an alternative independent evidence estimation tool. Then we propose using the NRE to predict the in concordance $R$ distribution, against which one can calibrate $R_{\rm obs}$. A schematic of the NRE or tensionnet is shown in fig. 1.
@@ VI Validating the NRE
$\mathcal{Z}=\mathcal{N}(m+M\mu,C+M\Sigma M^{\prime}).$ (20)
We use the lsbi package to evaluate these expressions  [^1] https://github.com/handley-lab/lsbi.
@@ VI.1 Assessing the performance of the NRE
For each $\Sigma$ we generate $500,000$ matched observations from experiment $A$ and experiment $B$ for training the NRE. We use an exponentially decaying learning rate with an initial value of $10^{-3}$, a step size of 1000 and a decay rate of 0.9. We use a ReLU activation function in the hidden layers, five hidden layers of 25 nodes each, a maximum number of epochs of 1000 with early stopping and a batch size of 1000. We use the ADAM optimizer for training. Once trained, we generate a new set of 5000 previously unseen in concordance observations from the models to put through the NRE and generate a predicted distribution of $\log R$.
The top panel of fig. 3 shows the predicted distributions (dashed lines) from the NRE versus the analytic distributions (solid lines) for different $\Sigma$. The solid black line shows the sigmoid activation function. The bottom three panels show the predicted versus true $\log R$ for each pair of data samples in the distribution. As the prior widens and $\log R$ becomes larger, the accuracy with which the distribution is recovered degrades as expected. Performance drops off, particularly for large prior widths, when $\log R>10$. Caution needs to therefore be taken when using the NRE to calibrate values of $\log R\gg 10$. In such circumstances, one could consider reducing the width of the prior or running nested sampling on a handful of simulations to gauge how well the NRE is performing. If all one is interested in is the tension between different data sets, one could also choose one’s prior so that $\log R$ is closer to 1 since the proposed tension metrics $T$ and $C$ are prior independent. For the orange distribution with $\Sigma=0.1\mathcal{I}$ and to some extent the purple distribution with $\Sigma=1\mathcal{I}$, the NRE accurately recovers the $\log R$ distribution.
Figure 3: To illustrate the performance of the tensionnet we hypothesise two experiments observing data that can be described with a linear model and a Gaussian likelihood function. By then defining our prior to also be Gaussian with a diagonal covariance $\Sigma$ we can analytically calculate the joint and individual evidences and the tension statistic $R$. We draw a test from the joint distribution $\mathcal{Z}(D_{A},D_{B})$ set which we use to analytically derive the in concordance $\log R$ distribution (solid lines, top panel) and predict the distribution from the NRE (dashed lines, top panel) for different prior widths. We also show the sigmoid activation function for reference. The bottom row shows the predicted versus true $\log R$ values for the test set for different prior widths. Performance begins to break down for $\log R>10$ however, for narrower priors corresponding to lower values of $\Sigma$ the tensionnet correctly recovers the in concordance $\log R$ distribution.
@@ VI.2 Calibrating out the prior
Using the above example, we can also illustrate how the tensionnet can be used to calibrate out the dependence of $R$ on the prior. In fig. 4, we keep our data model the same but change the prior width on our three parameters. Our observed dataset is drawn from the narrowest prior and kept the same throughout. We can clearly see that as the prior width increases, so does $R_{\mathrm{obs}}$ as expected. However, we can also see that the true distribution (purple) of in concordance $\log R$ values also shifts to higher values. When we use this distribution to calibrate $\log R_{\mathrm{obs}}$ into $T$ and $C$ the values are approximately constant regardless of the prior width. Calibrating against the predicted distribution from the NRE (orange) gives largely consistent results with some degradation in performance for the largest prior width as expected from the last section. We repeat the analysis five times and report the average values of $T$ and $C$ with an associated error for both the true and predicted distributions in fig. 4.
Figure 4: Using the linear model described in section VI we show how the in concordance $\log R$ distribution can be used to calibrate the prior dependence of the $R$ statistic. We also show how the predicted in concordance $R$ distribution from the tensionnet is largely consistent with the analytic distribution. The narrowest prior is on the top row and the widest on the bottom row. The first column shows the distribution of in concordance $\log R$ values calculated analytically in purple and as predicted by the NRE in orange. We also show the analytically calculated value of $\log R$ for a simulation drawn from the narrow prior as a red dashed line. The middle column shows the CDFs derived from the two in concordance distributions and as horizontal dashed lines the value of the CDF at $R_{\mathrm{obs}}$ according to the analytic (purple) and NRE (orange) distributions. The final column shows the average values over five runs of $T$ and $C$ derived using the true analytic distributions and the NRE for each prior with an associated error. From the first column of the figure, we see clearly see the prior dependence of the $\log R$ distribution. However, we can also see that the values of $T$ and $C$ predicted with the true in concordance $\log R$ distribution remain approximately constant regardless of the prior width. We can also see that the calibrating with the predicted and analytic distributions give largely consistent results.

### v2

@@ I Introduction
Independently confirming conclusions about the nature of our Universe from one experiment with another is crucial to the advance of knowledge. When the inference from two different experiments disagree with each other, we call this tension. Tension between different datasets raises questions about the need for new physics and better descriptions of our instruments and systematics. The $H_{0}$ and $\sigma_{8}$ tensions (e.g. knox2020HubbleHunters; Efstathiou2020LockdownH0; Amon2022s8; Preston2023s8; Abbott2023DESKIDS) are the most commonly encountered examples in cosmology, but other examples include observations of the 21-cm signal from cosmic dawn (Singh_SARAS3_2022), tensions in the amplitude of the matter power spectrum (Battye_tensions_2015) and tension in estimates of the curvature of the Universe Handley2019Curvature. A historical example of tension is in the measurement of the matter density $\Omega_{m}$ Peebles1984OmegaM; Krauss1995OmegaM; Ostriker1995OmegaM; Efstathiou1990OmegaM which was resolved by the discovery of the accelerating universe Riess1998Accelerating. A detailed review of cosmological tensions can be found in Abdalla2022TensionReview.
@@
Many different measures of tension have been proposed and are widely used in cosmological studies. Examples include Bayesian Suspiciousness Handley_tensions_2019, estimators of the probability of observed parameter differences Raveri2020parameterDifference; Raveri2021ParameterDiff, Goodness of fit degradation Raveri2019GoFDegredation and Eigentension Park2020Eigentension. These tension statistics are summarised and reviewed in Charnock_2017; Lemos2021DESTensions and Saraivanov2024Metrics. It is common practice to rephrase these tension metrics in $\sigma$ units of tension, corresponding to probabilities on a one dimensional normal distribution. When expressed in this way, one would expect that the various tension metrics predict the same level of tension or concordance between different datasets. However, this is often not the case due to the various assumptions that are made when defining the statistics. To tackle this issue, we can try to define tension metrics that do not make these assumptions, however, the tension statistics often lose some of their interpretability when we do this. Instead, we try to calibrate out these assumptions in sensible ways. Calibration of tension statistics is an important and often overlooked step that needs to be taken to correctly interrogate the tension between different datasets.
@@
A commonly used metric of tension is $R$ corresponding to the ratio of a joint evidence and the product of individual evidences for two different datasets under a common model. The $R$ statistic was first proposed in (Marshall_2006) and has been used to quantify tension in a number of cosmological studies (e.g. Trotta_2008; Seehars_2016). $R$ has also been used to perform model comparison in some works, although the authors of Cortes2024Tension showed that this approach to model comparison is incomplete.
@@
The ratio $R$ suffers from a non-trivial dependence on the prior that is not always accounted for properly (Handley_tensions_2019). In Handley_tensions_2019 the authors showed that as one decreases the prior width on the common parameters in the model of the two datasets, then the tension between the datasets should increase and $R$ should decrease. Intuitively, one can see that it is more satisfying if the two experiments favour parameters that are close together given a wide prior in comparison to a narrow prior. However, what constitutes ‘narrow’ and ‘wide’ is problem specific and subjective, making the interpretation of $R$ difficult. We would like to calibrate out the prior dependence.
[…]
@@
In Lemos2021DESTensions the authors demonstrate that simulations can be used to calibrate tension metrics with the Planck data and Dark Energy Survey (DES) data. They proposed taking a fiducial set of parameters, such as the maximum posterior point for Planck, shifting these parameter values by some posterior-informed step sizes to induce a known degree of tension, simulate the now in tension DES observations and calculate the value of ones chosen tension metric between the real Planck data and the simulation. To do this, however, one often has to run expensive sampling algorithms on the simulated data to calculate tension statistics such as $R$ and $S$.
@@
We propose calibrating the prior dependence of $R$ using neural ratio estimation (NRE) (e.g. Cranmer202SBI; Miller2021TMNRE; Cole2022TMNRE). NREs are classifiers, with interpretable outputs, that determine whether two quantities are drawn from independent distributions or a joint distribution. We show that the output of an NRE trained on simulations of two experiments observables can be interpreted as $R$ and that an appropriately trained NRE can be used to calibrate for the prior dependence of the dimensionless $R$ statistic. Since $R$ requires the calculation of three Bayesian evidences, it is an expensive statistic to evaluate using traditional methods like nested sampling. We show that $R$ can be accessed with a significantly smaller computational overhead using cutting edge machine learning tools. We call our NRE setup the tensionnet.
@@
In section˜II we summarise Bayesian inference and give more details about $R$. In section˜III we discuss the interpretation of $R$ and follow this with a discussion on NREs in section˜IV. We discuss calibrating $R$ with NREs in section˜V. We then test our method on toy examples with known in concordance $R$ distributions in section˜VI. We then apply the tensionnet to a toy example inspired by 21-cm cosmology and to assess the tension between Baryon Acoustic Oscillations (BAO) observations from the Dark Energy Spectroscopic Instrument (DESI) and the Sloan Digital Sky Survey (SDSS), in section˜VII. We consider some limitations of our method in section˜VIII and conclude in section˜IX.
@@ II Bayesian Inference and Tension Statistics
$\mathcal{Z}=\int\mathcal{L}(\theta)\pi(\theta)d\theta.$ (2)
An efficient and accurate way to recover both the evidence and the posterior is with nested sampling (skilling_ns_2006; Ashton_NS_2022) although other methods exist (e.g. Trotta_2007; Heavens_2017; Srinivasan2024floZ; emcee; Polanska2023harmonicmean).
@@
$R$ $=\frac{\mathcal{Z}_{A,B}}{\mathcal{Z}_{A}\mathcal{Z}_{B}}=\frac{1}{\mathcal{Z}_{A}\mathcal{Z}_{B}}\int\mathcal{L}_{A}\mathcal{L}_{B}\pi d\theta$ (4)
$=\int\frac{\mathcal{L}_{A}\pi}{\mathcal{Z}_{A}}\frac{\mathcal{L}_{B}\pi}{\mathcal{Z}_{B}}\frac{1}{\pi}d\theta=\int\frac{P_{A}P_{B}}{\pi}d\theta$
@@
where we have assumed the data sets are independent, and the angled brackets represent averages over the distributions $P_{A}$ and $P_{B}$ (Handley_tensions_2019). For a uniform prior, $\pi=1/V$ where $V$ is the volume, one can see from eq.˜4 that if the prior is made smaller than $R$ being proportional to $V$ also decreases. This logic generalises to more complicated priors.
@@ III Interpreting $R$
$R$ has several attractive properties Handley_tensions_2019;
  • Dimensionally consistent: Whilst the Bayesian evidence typically carries units that are the inverse of the measure on the data space, by taking the ratio of evidences the units cancel. $R$ is therefore not dependent on the units of the model space.
  • Parameterisation Invariant: The value of $R$ does not change if you reparameterise the model (e.g. switch from mass to log mass). While priors and likelihoods can change under reparameterisations the Bayesian evidence should properly account for this and ratios of evidences will remain invariant to these changes. $R$ therefore reflects the true model-data compatibility and is not sensitive to arbitrary coordinate choices.
  • Symmetric: Swapping the two data sets does not affect the value of $R$. So there is no preferred ordering and both data sets are treated equally.
It is typically interpreted with respect to a value of 1 with $R\ll 1$ corresponding to inconsistent datasets and $R\gg 1$ to consistent data. However, this interpretation does not tell us the degree to which our datasets are in tension given the prior and model choice. To try and quantify the tension between observations from the Dark Energy Survey and Planck, the authors of DES_Y1 interpreted $R$ on a Jefferys’ scale(jeffreys1983theory). The Jefferys’ scale is, however, somewhat arbitrary.
In Handley_tensions_2019 the authors showed that
@@
When interpreting $R$ one has to keep in mind the impact which the prior has on its value. Reducing the width of the prior will increase the apparent tension between the datasets by reducing the value of $R$. The authors of Handley_tensions_2019 suggest repeating our analysis with sensible modifications to the prior distribution to determine how stable the value of $R$ is and consequently our conclusions regarding the tension between different datasets.
@@ IV Neural Ratio Estimation
$D_{A}$ $\vdots\vdots$ $D_{B}$ $\vdots\vdots$ $\log R$ $p=S_{\sigma}(\log R)$ Loss Function: $l=\frac{1}{N}\bigg[\sum_{i}^{N}y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})\bigg]$ Neural Ratio Estimation of $\log R=\log\frac{P(D_{A},D_{B})}{P(D_{A})P(D_{B})}$
Figure 1: A schematic of the neural ratio estimator (NRE) used in this work, which we refer to as a tensionnet. The NRE is trained on matched and mismatched pairs of simulated observations from two different experiments AA and $B$ and outputs an estimate of the tension statistic $R$. The network is trained using the binary cross entropy loss function.
@@ V Calibrating $R$ with NREs
We propose that the true observed tension $R_{\mathrm{obs}}$ is calculated using nested sampling (e.g. Handley_tensions_2019) or an alternative independent evidence estimation tool. Then we propose using the NRE to predict the in concordance $R$ distribution, against which one can calibrate $R_{\rm obs}$. A schematic of the NRE or tensionnet is shown in fig.˜1.
@@ VI Validating the NRE
$\mathcal{Z}=\mathcal{N}(m+M\mu,C+M\Sigma M^{T}).$ (20)
where $M^{T}$ is the transpose of $M$. We use the lsbi package to evaluate these expressions  [^1] https://github.com/handley-lab/lsbi.
@@ VI.1 Assessing the performance of the NRE
For each $\Sigma$ we generate $500,000$ matched observations from experiment $A$ and experiment $B$ for training the NRE. We use an exponentially decaying learning rate with an initial value of $10^{-3}$, a step size of 1000 epochs and a decay rate of 0.9. We use a ReLU activation function in the hidden layers, five hidden layers of 25 nodes each, a maximum number of epochs of 1000 with early stopping and a batch size of 1000. We use the ADAM optimizer for training. Once trained, we generate a new set of 5000 previously unseen in concordance observations from the models to put through the NRE and generate a predicted distribution of $\log R$.
The top panel of fig.˜3 shows the predicted distributions (dashed lines) from the NRE versus the analytic distributions (solid lines) for different $\Sigma$. The solid black line shows the sigmoid activation function. The bottom three panels show the predicted versus true $\log R$ for each pair of data samples in the distribution. As the prior widens and $\log R$ becomes larger, the accuracy with which the distribution is recovered degrades as expected. Performance drops off, particularly for large prior widths, when $\log R>10$. Caution needs to therefore be taken when using the NRE to calibrate values of $\log R\gg 10$. In such circumstances, one could consider reducing the width of the prior or running nested sampling on a handful of simulations to gauge how well the NRE is performing. If all one is interested in is the tension between different data sets, one could also choose one’s prior so that $\log R$ is closer to 1 since the proposed tension metrics $T$ and $C$ are prior independent. For the orange distribution with $\Sigma=0.1\mathcal{I}$ and to some extent the purple distribution with $\Sigma=1\mathcal{I}$, the NRE accurately recovers the $\log R$ distribution.
Figure 3: We hypothesise two experiments observing data that can be described with a linear model and a Gaussian likelihood function. By then defining our prior to also be Gaussian with a diagonal covariance $\Sigma$ we can analytically calculate the joint and individual evidences and the tension statistic $R$. We draw a test set from the joint distribution $\mathcal{Z}(D_{A},D_{B})$ which we use to analytically derive the in concordance $\log R$ distribution (solid lines, top panel) and predict the distribution from the NRE (dashed lines, top panel) for different prior widths. We also show the sigmoid activation function for reference. The bottom row shows the predicted versus true $\log R$ values for the test set for different prior widths. Performance begins to break down for $\log R>10$.
@@ VI.2 Calibrating out the prior
Using the above example, we can also illustrate how the tensionnet can be used to calibrate out the dependence of $R$ on the prior. In fig.˜4, we keep our data model the same but change the prior width on our three parameters. Our observed dataset is drawn from the narrowest prior and kept the same throughout. We can clearly see that as the prior width increases, so does $R_{\mathrm{obs}}$ as expected. However, we can also see that the true distribution (purple) of in concordance $\log R$ values also shifts to higher values. When we use this distribution to calibrate $\log R_{\mathrm{obs}}$ into $T$ and $C$ the values are approximately constant regardless of the prior width. Calibrating against the predicted distribution from the NRE (orange) gives largely consistent results with some degradation in performance for the largest prior width as expected from the last section. We repeat the analysis five times and report the average values of $T$ and $C$ with an associated error for both the true and predicted distributions in fig.˜4.
Figure 4: Using the linear model described in section˜VI we show how the in concordance $\log R$ distribution can be used to calibrate the prior dependence of the $R$ statistic. We also show how the predicted in concordance $R$ distribution from the tensionnet is largely consistent with the analytic distribution. The narrowest prior is on the top row and the widest on the bottom row. The first column shows the distribution of in concordance $\log R$ values calculated analytically in purple and as predicted by the NRE in orange. We also show the analytically calculated value of $\log R$ for a simulation drawn from the narrow prior as a red dashed line. The middle column shows the CDFs derived from the two in concordance distributions and as horizontal dashed lines the value of the CDF at $R_{\mathrm{obs}}$ according to the analytic (purple) and NRE (orange) distributions. The final column shows the average values over five runs of $T$ and $C$ derived using the true analytic distributions and the NRE for each prior with an associated error.